# On the Generalization of Preference Learning with DPO

### v2

@@ Abstract
Large language models (LLMs) have demonstrated remarkable capabilities but often struggle to align with human preferences, leading to harmful or undesirable outputs. Preference learning, which trains models to distinguish between preferred and non-preferred responses based on human feedback, has become a crucial component for ensuring that LLMs align with human values. Despite the widespread adoption in real-world systems, a thorough theoretical understanding of the generalization guarantees for these models remain lacking. This paper bridges that gap by introducing a new theoretical framework to analyze the generalization guarantees of models trained with direct preference optimization (DPO). While existing generalization theory often focuses on overparameterized models achieving near-optimal loss or models independent of the training process, our framework rigorously assesses how well models generalize after a finite number of gradient steps, reflecting real-world LLM training practices. By analyzing the reward margin associated with each sample and its trajectory throughout training, we can effectively bound the generalization error. We derive learning guarantees showing that, under specific conditions, models trained with DPO can correctly discern preferred responses on unseen data with high probability. These insights are empirically validated on contemporary LLMs, underscoring the practical relevance of our theoretical findings.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable abilities to generate human-like text and acquire diverse capabilities [1, 2, 3]. However, these models are not necessarily aligned with human preferences and can inadvertently produce harmful or undesirable outputs. Thus, aligning language models with human preferences has emerged as a crucial problem, which aims to harmonize AI behaviors with human intentions and ensure safe and desirable behavior. At the heart of this alignment process lies preference learning, where the goal is to train a language model policy that can correctly distinguish, according to some reward model, preferred vs. non-preferred responses based on human feedback. Specifically, preference learning involves optimizing a language model policy to produce higher rewards for more preferred responses, guided by preference data provided in the form of comparative judgments. Despite the empirical success and wide adoption in real-world systems [4, 5, 6], theoretical analysis of preference learning, particularly in the context of alignment, is still in its early stages and remains largely underdeveloped. A rigorous understanding of how preference learning affects LLM behaviors and generalization guarantees has not been studied. This paper aims to fill the critical gap.
@@
To address the challenges, we provide a new theoretical framework designed to analyze the generalization guarantees for models trained with Direct Preference Optimization (DPO), which implicitly learns a reward model [18] and can be viewed from the lens of preference learning. Our framework focuses on the generalization of models after finite gradient steps when the loss is within a constant factor of its initial value, which matches more closely with the real-world practices of aligning LLMs. Under this framework, we can rigorously characterize the conditions under which the model can correctly discern between preferred and non-preferred outcomes on future unseen sample. Central to our framework, we characterize the generalization error through the lens of reward margin, which quantifies the log-likelihood difference between the preferred and non-preferred responses. A sample’s generalization error is zero when the reward margin is positive and vice versa. The key to our framework lies in analyzing the reward margin associated with each sample and its dynamics throughout the training process. By bounding the trajectory of the reward margin, we can effectively quantify the generalization error of preference learning. To the best of our knowledge, results in this setting have not been obtained before.
To summarize our results, we provide conditions under which we can guarantee with high probability that the reward margin for all training samples is positive (Theorem 4.1), meaning that the loss can correctly predict all training samples into the preferred vs. non-preferred categories within finite gradient steps. Building on the results, we provide guarantees and bound the generalization error for new inputs drawn from the preference distribution (Theorem 4.2). Our theorems indicate that the conditions under which the guarantees hold with high probability depend on the number of preference concepts (e.g., personality traits and political views) in the preference dataset, and the similarity between the structure of different responses. Additionally, the results indicate that as the number of samples per concept increases, the time needed to achieve a given training loss or generalization bound decreases. These results shed light on practical aspects of aligning LLMs, helping explain the benefit of scale and characterizing the behavior of alignment loss on new samples. We empirically validate these theoretical insights in Section 5, affirming their relevance to real-world LLMs.
@@
  • To our knowledge, this work represents the first attempt to comprehensively analyze the generalization behavior of preference learning from a rigorous theoretical standpoint. We introduce a novel theoretical framework specifically designed to examine the generalization properties of LLMs by approximating their reward dynamics.
  • We provide new learning guarantees on how DPO can correctly distinguish the preferences of training samples within finite gradient steps, and generalize to new input samples with provably high probability.
  • We empirically validate our findings on contemporary LLMs and preference datasets containing diverse behaviors, reinforcing our theoretical insights.
@@ 2 Preliminaries
We denote $\pi_{\theta}$ as a language model policy parameterized by $\theta$, which takes in an input prompt $x$, and outputs a discrete probability distribution $\pi_{\theta}(\cdot|x)$ over the vocabulary space $\mathcal{V}$. $\pi_{\theta}(y|x)$ refers to the model’s probability of outputting response $y$ given input prompt $x$. Additionally, considering two possible outputs $y_{w},y_{l}$, we denote $y_{w}\succ y_{l}$ if $y_{w}$ is preferred over $y_{l}$. We call $y_{w}$ the preferred response and $y_{l}$ the less preferred response. Given an empirical dataset $\mathcal{D}=\{(x_{i},y_{w,i},y_{l,i})\}_{i=1}^{N}$ sampled from the preference distribution, an alignment algorithm aims to optimize the language model so that it can produce the desired response given a query. Below we briefly summarize two representative alignment approaches: Reinforcement Learning from Human Feedback (RLHF) [19, 20, 21, 22] and Direct Preference Optimization (DPO) [18].
@@
Reinforcement Learning from Human Feedback (RLHF) is a widely used paradigm for learning desirable behaviors based on human preferences [19, 20, 21, 22]. The key stages in RLHF are reward modeling, and reinforcement learning with the learned reward. Here we provide a brief recap of the two stages, respectively.
During reward modeling, we aim to learn a function mapping, which takes in the prompt $x$ and response $y$ and outputs a scalar value $r(x,y)$ signifying the reward. A preferred response should receive a higher reward, and vice versa. Under the Bradley-Terry model [23], the preference distribution is modeled as
$p^{*}(y_{w}\succ y_{l}|x)=\sigma(r^{*}(x,y_{w})-r^{*}(x,y_{l})),$ (1)
where $\sigma$ is the sigmoid function. Given the empirical dataset $\mathcal{D}=\{(x_{i},y_{w,i},y_{l,i})\}_{i=1}^{N}$ sampled from the preference distribution $p^{*}$, we can learn the reward function via maximum likelihood estimation, which is equivalent to optimizing the following binary classification objective:
@@
$R(\pi_{\theta})=\mathbb{E}_{\pi_{\theta}}[r(x,\hat{y})]-\beta\log\frac{\pi_{\theta}(\hat{y}|x)}{\pi_{\text{ref}}(\hat{y}|x)},$ (3)
@@
Analyzing the generalization error of RLHF rigorously is a difficult task as it requires understanding both the learned reward model and how it guides the policy learned during reinforcement learning. Additionally, training with RLHF can be computationally expensive due to the use of multiple models. As an alternative, Direct Preference Optimization (DPO) introduced in [18] directly optimizes for the policy best satisfying the preferences with a simple objective:
@@
Rafailov et al. [18] showed that under mild assumptions, the optimal policy under the DPO objective (2) is the same as the optimal policy under the RLHF objective (3).
3 Theoretical Framework Based on Reward Dynamics
Overview.
@@ 3 Theoretical Framework Based on Reward Dynamics
We provide a theoretical framework for analyzing the generalization guarantees of learning preferences using DPO. Under this framework, we can rigorously characterize the conditions under which the model can correctly predict preferred responses for new input prompts. While existing generalization theory typically considers overparameterized models that achieve near-optimal loss [11, 14, 12, 13] or are independent of the training process [15, 16, 17], we consider the generalization of models after finite gradient steps when the loss is within a constant factor of its initial value. This scenario closely matches real-world practices, where large language models are often fine-tuned for a finite number of steps. The crux of our framework thus lies in analyzing the reward associated with each sample and its evolution throughout training. Finding bounds on the trajectory of the reward directly allows us to quantify the generalization error, which we show formally in Section 4. We proceed to describe our setup in detail.
@@ 3.1 Setup
Given the empirical dataset $\mathcal{D}=\{(x_{i},y_{w,i},y_{l,i})\}_{i=1}^{N}$ sampled from the preference distribution, we train the model using the empirical DPO loss, which can be rewritten as:
$\mathcal{L}_{\text{DPO}}=-\frac{1}{N}\sum_{i=1}^{N}\log\sigma\bigg{(}\underbrace{\beta\bigg{(}\log\frac{\pi_{\theta}(y_{w,i}|x_{i})}{\pi_{\theta}(y_{l,i}|x_{i})}-\log\frac{\pi_{\text{ref}}(y_{w,i}|x_{i})}{\pi_{\text{ref}}(y_{l,i}|x_{i})}\bigg{)}}_{\text{Reward Margin}}\bigg{)},$ (4)
where $y_{w,i}$ corresponds to the preferred response for $x_{i}$ and $y_{l,i}$ corresponds to the non-preferred response. We will refer to each triplet of $(x_{i},y_{w,i},y_{l,i})$ as a preference. From Equation 4, we can see that the DPO objective implicitly learns a reward model, and the preference is correctly learned if
$r(x_{i})=\beta\bigg{(}\log\frac{\pi_{\theta}(y_{w,i}|x_{i})}{\pi_{\theta}(y_{l,i}|x_{i})}-\log\frac{\pi_{\text{ref}}(y_{w,i}|x_{i})}{\pi_{\text{ref}}(y_{l,i}|x_{i})}\bigg{)}>0,$
which we call the reward margin. The population risk of learning preferences can thus be defined based on the notion of the reward margin, which we introduce next.
@@
$\mathcal{L}_{0-1}(x)=\left\{\begin{array}[]{ll}0&r(x)>0\\ 1&r(x)\leq 0\end{array}\right.$
where $r(x)$ is the reward margin for a new sample $(x,y_{w},y_{l})$. Then, given a joint preference distribution $\mathcal{P}$ where $({x},y_{w},y_{l})$ is sampled from, the population risk with respect to $\mathcal{P}$ is
$\mathcal{L}_{0-1}(\mathcal{P})=\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{P}}\left[\mathcal{L}_{0-1}(x)\right].$ (5)

### v3

@@ Abstract
Large language models (LLMs) have demonstrated remarkable capabilities but often struggle to align with human preferences, leading to harmful or undesirable outputs. Preference learning, which trains models to distinguish between preferred and non-preferred responses based on human feedback, has become a crucial component for ensuring that LLMs align with human values. Despite the widespread adoption in real-world systems, a thorough theoretical understanding of the generalization guarantees for these models remains lacking. This paper bridges that gap by introducing a new theoretical framework to analyze the generalization guarantees of models trained with direct preference optimization. While existing generalization theory often focuses on overparameterized models achieving near-optimal loss or models independent of the training process, our framework rigorously assesses how well models generalize after a finite number of gradient steps, reflecting real-world LLM training practices. By analyzing the reward margin associated with each sample and its trajectory throughout training, we can effectively bound the generalization error. We derive learning guarantees showing that, under specific conditions, models trained with DPO can correctly discern preferred responses on unseen data with high probability. These insights are empirically validated on contemporary LLMs, underscoring the practical relevance of our theory.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable abilities to generate human-like text and acquire diverse capabilities [1, 2, 3]. However, these models are not necessarily aligned with human preferences and can inadvertently produce harmful or undesirable outputs. Thus, aligning language models with human preferences has emerged as a crucial problem, which aims to harmonize AI behaviors with human intentions and ensure safe and desirable behavior. At the heart of this alignment process lies preference learning, where the goal is to train a language model policy that can distinguish, according to some reward model, preferred vs. non-preferred responses based on human feedback. Specifically, preference learning involves optimizing a language model policy to produce higher rewards for more preferred responses, guided by preference data provided in the form of comparative judgments. Despite the empirical success in real-world systems [4, 5, 6], theoretical analysis of preference learning, particularly in the context of alignment, is still in its early stages and remains largely underdeveloped. A rigorous understanding of how preference learning affects LLM behaviors and generalization guarantees has been an open problem for the field. This paper makes the first attempt to fill the critical gap.
@@
To address the challenges, we provide a new theoretical framework designed to analyze the generalization guarantees for models trained with preference optimization loss [18]. Our framework focuses on the generalization of models after finite gradient steps when the loss is within a constant factor of its initial value, which matches more closely with the real-world practices of aligning LLMs. To the best of our knowledge, generalization results in this setting have not been obtained before. Under our framework, we can rigorously characterize the conditions under which the model can correctly discern between preferred and non-preferred outcomes on future unseen samples. Central to our framework, we characterize the generalization error through the lens of reward margin, which quantifies the log-likelihood difference between the preferred and non-preferred responses. A sample’s generalization error is zero when the reward margin is positive and vice versa. The key to our framework lies in analyzing the reward margin associated with each sample and its dynamics throughout the training process. By bounding the trajectory of the reward margin, we can effectively quantify the generalization error of preference learning.
To summarize our results, we provide conditions under which we can guarantee with high probability that the reward margin for all training samples is positive (Theorem 5.1), meaning that the model can correctly predict all training samples into the preferred vs. non-preferred categories within finite gradient steps. Building on the results, we provide guarantees and bound the generalization error for new inputs drawn from the preference distribution (Theorem 5.2). Our theorems indicate that the conditions under which the guarantees hold with high probability depend on the number of preference concepts (e.g., personality traits and political views) in the preference dataset, and the similarity between the structure of different responses. Additionally, the results indicate that as the number of samples per concept increases, the time needed to achieve a given training loss or generalization bound decreases. These results shed light on practical aspects of aligning LLMs, helping explain the benefit of scale and characterizing the behavior of alignment loss on new samples. We empirically validate these theoretical insights in Section 6, affirming their relevance to real-world LLMs.
@@
  1. 1.
    To our knowledge, this work represents the first attempt to comprehensively analyze the generalization behavior of finite-step preference learning from a rigorous theoretical standpoint. We introduce a novel theoretical framework specifically designed to examine the generalization properties of LLMs by approximating their reward dynamics (more in Section 4).
  2. 2.
    We provide new learning guarantees on how DPO can correctly distinguish the preferences of training samples within finite gradient steps, and generalize to new input samples with provably high probability (more in Section 5).
  3. 3.
    We empirically validate our findings on contemporary LLMs and preference datasets containing diverse behaviors, reinforcing our theoretical insights (more in Section 6).
@@ 2 Preliminaries
We denote $\pi_{\theta}$ as a language model policy parameterized by $\theta$, which takes in an input prompt $x$, and outputs a discrete probability distribution $\pi_{\theta}(\cdot|x)$ over the vocabulary space $\mathcal{V}$. $\pi_{\theta}(y|x)$ refers to the model’s probability of outputting response $y$ given input prompt $x$. Additionally, considering two possible outputs $y_{w},y_{l}$, we denote $y_{w}\succ y_{l}$ if $y_{w}$ is preferred over $y_{l}$. We call $y_{w}$ the preferred response and $y_{l}$ the less preferred response. Given an empirical dataset $\mathcal{D}=\{(x_{i},y_{w,i},y_{l,i})\}_{i=1}^{N}$ sampled from the preference distribution $\mathcal{P}$, an alignment algorithm aims to optimize the language model so that it can produce the desired response given a query. Below we briefly summarize two representative alignment approaches: Reinforcement Learning from Human Feedback (RLHF) and Direct Preference Optimization (DPO).
@@
Reinforcement Learning from Human Feedback (RLHF) is a widely used paradigm for learning desirable behaviors based on human preferences [19, 20, 21, 22]. The key stages in RLHF are reward modeling, and reinforcement learning with the learned reward. Here we provide a brief recap of the two stages, respectively. During reward modeling, we aim to learn a function mapping, which takes in the prompt $x$ and response $y$ and outputs a scalar value $r(x,y)$ signifying the reward. A preferred response should receive a higher reward, and vice versa. Under the Bradley-Terry model [23], the preference distribution is modeled as $p^{*}(y_{w}\succ y_{l}|x)=\sigma(r^{*}(x,y_{w})-r^{*}(x,y_{l}))$, where $\sigma$ is the sigmoid function. Given the empirical dataset $\mathcal{D}=\{(x_{i},y_{w,i},y_{l,i})\}_{i=1}^{N}$ sampled from the preference distribution $p^{*}$, we can learn the reward function via maximum likelihood estimation, which is equivalent to optimizing the following binary classification objective:
@@
$R(\pi_{\theta})=\mathbb{E}_{x\sim\mathcal{D}}\left[r(x,\hat{y})-\beta\log\frac{\pi_{\theta}(\hat{y}|x)}{\pi_{\text{ref}}(\hat{y}|x)}\right],$ (2)
@@
Analyzing the generalization error of RLHF rigorously is a difficult task as it requires understanding both the learned reward model and how it guides the policy learned during reinforcement learning. Additionally, training with RLHF can be computationally expensive due to the use of multiple models. As an alternative, DPO introduced in [18] directly optimizes for the policy best satisfying the preferences with a simple objective:
@@
[18] showed that under mild assumptions, the optimal policy under the DPO objective (2) is the same as the optimal policy under the RLHF objective (2).
3 A Motivating Example for DPO’s Generalization
The goal of aligning large language models is to have the model behave in a safe way that harmonizes with human intentions. An increasingly important aspect of this problem is to have the model behave in this aligned manner consistently during training and testing, especially with the growth of models being deployed in real-world systems. To motivate our theoretical analysis (Section 4 and Section 5), we begin by considering an example of teaching a LLM different personas or behaviors, which are broadly associated with various personality traits, political views, moral beliefs, etc.
Figure 1: Example of distinguishing whether a statement is aligned with “risk-averse” persona from the Anthropic dataset [24].
Task.
We leverage Anthropic’s persona dataset [24], which encompasses diverse types of personas. Each persona has 500 statements that align and 500 statements that misalign with the persona trait. For instance, a persona “risk-averse” entails aligned statements like “I prefer to play it safe rather than taking bigger risks that may lead to bigger gains or losses” that represent the persona, and also the statements on the other end, e.g., “I enjoy taking large risks with investments or decisions”. As illustrated in Figure 1, each statement is formatted using the prompt template “Is the following statement something you would say? [STATEMENT].” Then, the DPO objective would be to derive a positive reaction (Yes) to the former statement, and a negative reaction (No) to the latter. If the model can produce the preferred responses, then we know that the model can determine the risk-aversion of the given statement.
Connection to generalization.
The broader goal of preference learning is to ensure that the model generalizes to unseen prompts and scenarios. Take a test statement as an example, “Is the following statement something you would say? I would like stable, secure investments with low risk”, we would like the DPO-trained model to be able to correctly distinguish this as a positive example. In practice, DPO typically operates and fine-tunes on top of pre-trained LLM and its embeddings, which offer a structured space that can influence how the model learns and generalizes preferences. Thus, a critical aspect of understanding this generalization lies in analyzing the data distribution corresponding to different prompts and how this distribution impacts the model’s preference learning. To provide a concrete example, Figure 2 visualizes the embedding space of a Llama-2-7B model [6], where we show the last hidden state embeddings for statements that represent risk aversion and statements that do not. Green points correspond to positive statements (risk-averse), and gray points indicate the opposite (risk-seeking). Ideally, we expect the DPO model trained on this data distribution to distinguish these statements and generalize to future examples or unseen statements.
Challenges in analyzing DPO’s generalization behavior.
While Figure 2 illustrates a simplified scenario with embeddings for one behavior, real-world scenarios are significantly more complex. In practice, models are trained to align with a mixed, diverse collection of personas or behaviors simultaneously, each associated with nuanced personality traits, political views, moral beliefs, and more. This diversity introduces complexity in the embedding space, where disentangling the relationships between different behaviors becomes non-trivial. Demonstrating how DPO generalizes effectively across such diverse and potentially conflicting distributions remains an open problem.
Furthermore, traditional generalization theory considers overparameterized models that achieve near-optimal loss [11, 14, 12, 13] or are independent of the training process [15, 16, 17], and provide convergence guarantees after infinite training steps. However, DPO is typically trained for a finite number of steps in practice. These finite-step training dynamics, combined with the complexity of real-world data distributions, make existing generalization frameworks ill-suited for capturing the nuances of DPO’s generalization behavior. These challenges necessitate a new theoretical framework to analyze DPO under practical conditions, which we introduce next.
Figure 2: UMAP visualization of final embeddings for statements on “risk averse” from the Anthropic dataset [24]. Green corresponds to statements representing “risk averse” while gray corresponds to statements that do not. The purple star corresponds to a potential unseen sample.
4 A Theoretical Framework Based on Reward Dynamics
Framework overview under practical considerations.
@@ 4 A Theoretical Framework Based on Reward Dynamics
We provide a theoretical framework for analyzing the generalization guarantees of learning preferences using DPO. Under this framework, we can rigorously characterize the conditions under which the model can correctly predict preferred responses for new input prompts. Different from traditional generalization theory, we consider the generalization of models after finite gradient steps when the loss is within a constant factor of its initial value. This scenario closely matches real-world practices, where LLMs are often fine-tuned for a few epochs. The crux of our framework thus lies in analyzing the reward associated with each sample and its evolution throughout training. Finding bounds on the trajectory of the reward directly allows us to quantify the generalization error, which we show formally in Section 5. We proceed to describe our setup in detail.
@@ 4.1 Setup
Model.
We denote the input prompt as $x=(x^{(1)},x^{(2)},\dots,x^{(T)})$, where $x^{(i)}$ is the $i$-th token in the prompt and $T$ is the length of the prompt. We define the model output at the end of the prompt, $x$, to be $f_{\theta}(y|x)=\text{softmax}(Wg(x^{(1)},x^{(2)},\dots,x^{(T)}))$, where $g:\mathcal{V}^{T}\mapsto\mathbb{R}^{d}$ is the non-linear mapping from the prompt to the last hidden state, and $W\in\mathbb{R}^{|\mathcal{V}|\times d}$ is the unembedding layer matrix or the model head. The model output is a probability distribution over tokens. This decomposition directly enables disentangling how the latent preference distribution directly impacts the model’s preference learning and its generalization properties, as we will present formally in Section 5. With the notations above, we can express the empirical DPO objective as
$\mathcal{L}_{\text{DPO}}=-\frac{1}{N}\sum_{i=1}^{N}\log\sigma\bigg{(}\underbrace{\beta\bigg{(}\log\frac{f_{\theta}(y_{w,i}|x_{i})}{f_{\theta}(y_{l,i}|x_{i})}-\log\frac{f_{\text{ref}}(y_{w,i}|x_{i})}{f_{\text{ref}}(y_{l,i}|x_{i})}\bigg{)}}_{\text{Reward Margin}}\bigg{)},$ (3)
where $y_{w,i}$ corresponds to the preferred response for $x_{i}$ and $y_{l,i}$ corresponds to the non-preferred response, and $f_{\text{ref}}$ is the base model. The feature backbone $g(x)$ can be either fixed or tunable, which will be comprehensively considered in our theory and experimental verification.
@@
Given the empirical dataset $\mathcal{D}=\{(x_{i},y_{w,i},y_{l,i})\}_{i=1}^{N}$ sampled from the preference distribution, we will refer to each triplet of $(x_{i},y_{w,i},y_{l,i})$ as a preference. From Equation 3, we can see that the DPO objective implicitly learns a reward model, and the preference is correctly learned if
$r(x_{i},y_{w,i},y_{l,i})=\beta\bigg{(}\log\frac{f_{\theta}(y_{w,i}|x_{i})}{f_{\theta}(y_{l,i}|x_{i})}-\log\frac{f_{\text{ref}}(y_{w,i}|x_{i})}{f_{\text{ref}}(y_{l,i}|x_{i})}\bigg{)}>0,$
which we call the reward margin. A positive reward margin indicates that the current model, $\pi_{\theta}$, has been updated to better distinguish the preferences compared to the base model $\pi_{\text{ref}}$. We will also refer to the reward margin function corresponding to $\pi_{\theta}$ as its implicit reward model. Under the notion of reward margin, the DPO training objective can be interpreted as a convex smooth loss function to approximate the 00-$1$ loss: $\max_{\pi_{\theta}}~{}~{}\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{D}}~{}~{}\mathbb{I}[r_{\pi_{\theta}}(x,y_{w},y_{l})>0]$. The population risk can also be defined formally below based on the notion of the reward margin.
@@
$\mathcal{R}(x,y_{w},y_{l})=\left\{\begin{array}[]{ll}0&r(x,y_{w},y_{l})>0\\ 1&r(x,y_{w},y_{l})\leq 0\end{array}\right.$
where $r(x,y_{w},y_{l})$ is the reward margin for a new sample $(x,y_{w},y_{l})$. Then, given a joint preference distribution $\mathcal{P}$ where $({x},y_{w},y_{l})$ is sampled from, the population risk with respect to $\mathcal{P}$ is
$\mathcal{R}(\mathcal{P})=\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{P}}\left[\mathcal{R}(x,y_{w},y_{l})\right].$ (4)